OpenStudy (anonymous):

Help !!:/ Raise the number to the indicated power and express in trigonometric notation [2(cos118degree+ i sin 118 degree)]^4

OpenStudy (anonymous):

much easier than it looks take \(2^4=16\) and then multiply each angle by \(4\) i.e. \(4\times 118\)

OpenStudy (anonymous):

more explain plz

OpenStudy (anonymous):

e^(i x)=cos(x)+i sin(x)

OpenStudy (anonymous):

(cos(x)+i sin(x))^n=(e^(i x))^n=e^(i n x)=cos(n x)+i sin(n x)

OpenStudy (unklerhaukus):

note that eashy's formulae are in radians

OpenStudy (anonymous):

help

OpenStudy (unklerhaukus):

\[[2(\cos118°+ i \sin 118°)]^4\\ =[2(\cos(118°\tfrac\pi{180°})+ i\sin (118°\tfrac\pi{180°}))]^4\\ =[2e^{i(118°\tfrac\pi{180°})}]^4\\=2^4e^{4i(118°\tfrac\pi{180°})}\\ =2^4(\cos(4\times118°\tfrac\pi{180°})+ i\sin (4\times 118°\tfrac\pi{180°}))\]